(a)The table below show Business statistics I examination marks for a class of one hundred students in Mount Kenya university.

Marks

10-19

20-29

30-39

40-49

50-59

60-69

Number of students

4

8

a

22

48

b

The mean score of the students was 46.5

         (i) Show that the values of a and b are 12 and 6 respectively.                       (5marks)



1
Expert's answer
2021-10-22T11:59:40-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Marks & Number\ of \ students & C.F. \\ \hline
 10-19 & 4 & 185 \\
 \hdashline
 20-29 & 8 & 12 \\
 \hdashline
 30-39 & a & 12+a \\
 \hdashline
 40-49 & 22 & 34+a \\
 \hdashline
 50-59 & 48 & 82+a \\
 \hdashline
 60-69 & b & 82+a+b \\
 \hdashline
\end{array}"

The total frequency is "82+a+b"


"82+a+b=100"

"b=18-a"

The mean score of the students was 46.5.


"14.5(4)+24.5(8)+34.5a+44.5(22)"

"+54.5(48)+64.5b=3849+34.5a+64.5b"

"mean=\dfrac{3849+34.5a+64.5b}{82+a+b}=46.5"




"3849+34.5a+64.5b=3813+46.5a+46.5b"

"18b+36=12a"

"3b=2a-6"

Substitute


"3(18-a)=2a-6"

"5a=60"

"a=12"

"b=6"